# VIRTUE: Visual-Interactive Text-Image Universal Embedder

### v1

@@ Abstract
Multimodal representation learning models have demonstrated successful operation across complex tasks, and the integration of vision-language models (VLMs) has further enabled embedding models with instruction-following capabilities. However, existing embedding models lack visual-interactive capabilities to specify regions of interest from users (e.g., point, bounding box, mask), which have been explored in generative models to broaden their human-interactive applicability. Equipping embedding models with visual interactions not only would unlock new applications with localized grounding of user intent, which remains unexplored, but also enable the models to learn entity-level information within images to complement their global representations for conventional embedding tasks. In this paper, we propose a novel Visual-InteRactive Text-Image Universal Embedder (VIRTUE) that extends the capabilities of the segmentation model and the vision-language model to the realm of representation learning. In VIRTUE, the segmentation model can process visual prompts that pinpoint specific regions within an image, thereby enabling the embedder to handle complex and ambiguous scenarios more precisely. To evaluate the visual-interaction ability of VIRTUE, we introduce a large-scale Segmentation-and-Scene Caption Retrieval (SCaR) benchmark comprising 1M samples that aims to retrieve the text caption by jointly considering the entity with a specific object and image scene. VIRTUE consistently achieves a state-of-the-art performance with significant improvements across 36 universal MMEB (3.1%–8.5%) and five visual-interactive SCaR (15.2%–20.3%) tasks111The code, model, and benchmark will be released upon acceptance..
@@ 1 Introduction
Although VLM-based embedding models support interactive use, they only accept text as the human-machine interaction modality. In contrast, visual prompts, which have recently attracted attention in generative applications (You et al., 2024; Yuan et al., 2024; Lian et al., 2025), serve as an important but overlooked interaction channel. Visual prompting can not only enhance the downstream generation performance (Li et al., 2022b) but also provide precise spatial localization for fine-grained understanding (Liu et al., 2024c). This is particularly advantageous for embedding-based tasks, as it allows models to respond to visual inputs from the user beyond traditional global matching by capturing entity-level cues, thereby improving retrieval precision and alignment while complementing global representations.
@@
Method Novelty: We propose VIRTUE, a visual-interactive text-image universal embedder consisting of a VLM as well as a segmentation model to enable the visual interaction modality for human interactions. The segmentation model allows users to optionally provide different types of visual prompts via its prompt encoder and reinforces VIRTUE to capture entity-level representations in addition to global context.
@@ 2 Related Works
Multimodal Representation Learning. Early progress in text-image representation learning was driven by two-tower contrastive models that learn a joint embedding space by aligning an image encoder with a text encoder (e.g., CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), SigLIP (Zhai et al., 2023), OpenCLIP (Cherti et al., 2023)). These models provide effective global image-text matching and have served as foundation models for building vision-language models (VLMs) that tackle zero-shot downstream tasks (Tong et al., 2024; Wang et al., 2025a; Deitke et al., 2025). Subsequent work has advanced embedding performance along various dimensions. For instance, UniIR (Wei et al., 2024) finetunes CLIP/BLIP with the late fusion of text and image embeddings, and UniME (Gu et al., 2025) distills text knowledge from an LLM followed by two-stage negatives contrastive learning. Magiclens (Zhang et al., 2024) incorporates open-ended instructions into dual-encoder architectures with training on large-scale instruction datasets. More recently, VLMs that accept arbitrary mixtures of visual and textual inputs and are trained with instruction-style objectives have emerged as flexible and unified embedding providers that better perform on multimodal reasoning and compositional queries, e.g., E5-V (Jiang et al., 2024), VLM2Vec (Jiang et al., 2025b), GME (Zhang et al., 2025), and LamRA (Liu et al., 2025b). Despite these advancements, existing embedding models only support textual instructions and lack native support for direct visual prompts; they are typically trained only on holistic image-text alignment (Oh et al., 2024). In contrast, VIRTUE integrates a segmentation model with a pretrained VLM to fuse segmentation-derived, prompt-conditioned entity representations with global representations, producing unified embeddings that are both entity-aware and context-preserving, as evaluated on both MMEB and our newly constructed SCaR.
@@ 3.1 SCaR Overview
Task Definition. The main difference between conventional image-to-text retrieval (e.g., MSCOCO_i2t in MMEB) and visual-interactive image-to-text retrieval is the additional region-of-interest input. Formally, given an input image and a bounding box with ten candidates , the goal is to find the most relevant text caption , where is the embedding model and sim denotes cosine similarity. Since SCaR requires models to reason about both the specified object and the broader scene, candidate captions are intentionally challenging and demand joint reasoning over fine-grained bounding-box details and global image context.
@@ 3.2 Collection Pipeline
Fig. 2 illustrates the collection pipeline to construct SCaR, where all datasets are first converted to the COCO format to enable unified processing. To balance evaluation efficiency and coverage, we randomly sample up to five objects per image. For each sample, the image size, original caption, object category, and bounding box222While some datasets provide segmentation masks, we observed that GPT-4V struggles to interpret them reliably. In contrast, bounding boxes are easier for GPT-4V to parse and align with prior prompt-based collection strategies (Wang et al., 2025b; Jiang et al., 2025a); therefore, we opt for bounding boxes for SCaR. are provided in the prompt template, as shown in Fig. 4. The prompt instructs GPT-4V to return a JSON object containing the ground-truth caption and nine negatives. Since some datasets lack complete descriptions in terms of <object> <relation> <scene> (e.g., ADE20k only provides object names, and some RefCOCO+ samples do not contain global scene context), the prompt asks GPT-4V to verify whether the caption contains all three elements, and to supplement any missing elements with careful and image-grounded descriptions.
@@ 6 Conclusion
The codebase and partial SCaR benchmark samples (due to file size) are provided in the supplementary materials for reproducibility, and the experimental details as well as model configurations are summarized in Sec. 5.1 and Appendix C. VIRTUE-2B and VIRTUE-7B will be released in the camera-ready paper to advance the community with the realm of interactive representation learning. Evaluations on the MMEB benchmark follows their official use, while SCaR is detailed in Sec. 3 and Appendix D.
@@ D.4 Additional SCaR Statistics
We further analyze the composition of SCaR to better understand its benchmark characteristics. Fig. 6 summarizes detailed statistics, including word clouds, annotation counts, and image distributions for both the training and evaluation splits. Attributed to the creative and diverse instruction collection process described in Sec. 3.2, we observe that the word frequency distributions of ground-truth and negative captions are closely aligned across splits, suggesting that the benchmark is not dominated by a small set of frequent words. Moreover, both image- and annotation-level distributions reveal that COCO-Stuff and VisualGenome contribute the largest proportions of samples, reflecting their broad coverage of scenes and object relationships. Fig. 7 illustrates the sentence length distributions of ground-truth and negative captions across the five datasets in SCaR. Overall, the two distributions are closely aligned, indicating that the synthesized negative captions generated by GPT-4V preserve similar linguistic complexity to the ground-truth annotations rather than introducing trivial artifacts. We also observe slightly heavier tails in the negative caption distributions, suggesting greater variability in sentence length. This property enhances the linguistic diversity within SCaR while maintaining a comparable level of difficulty, thereby preventing models from exploiting superficial cues such as caption length.
@@ D.5 SCaR Examples
| Dataset | Query Image with bbox | Candidates | ||||||||||
@@
Appendix E Extensive Experiments
@@ E.1 Ablation Study
We conduct seven ablation variants to investigate the relative contributions of different components: 1) VLM backbones, 2) with and without an instruction in the query, 3) input image resolution, 4) segmentation streamline alternatives, 5) different sizes of SAM-2, and 6) different numbers of MLPs in the segmentation-language connector (Eq. 1), where one MLP projects from to , and three MLPs project from to 768, then to 1024, and finally to ; and 7) segmentation embedding lengths . For 4) segmentation streamline alternatives, we replace our segmentation model with Prompter from CLOC555CLOC cannot be directly evaluated or reproduced, as its model, code, and parts of training data are not publicly available. We attempted to contact the authors but did not receive a response. (Chen et al., 2025), which encodes bounding boxes via a randomly initialized Transformer, or directly crops images for the vision encoder (denoted as Cropped). Both variants concatenate embeddings from nine random crops, similar to the setting of nine sampled points in SAM2.1 for the segmentation-language connector for MMEB.
As shown in Tab. 6, although Phi-3.5-V has more parameters than Qwen2-VL-2B, the latter yields better results on MMEB and SCaR. Removing text instructions and using lower resolutions lead to a significantly inferior performance. The degraded results from Prompter and cropping suggest that pre-trained SAM2.1 provides not only more precise segmentations than simple crops but also a stronger understanding from its pre-trained knowledge. Although SAM2.1-L performs on par with SAM2.1-B+ in terms of MMEB, its performance degrades on SCaR. In contrast, using SAM2.1-S deleteriously impacts both benchmarks. Additionally, increasing MLP layers and segmentation embedding length does not consistently yield gains across MMEB and SCaR. While increasing segmentation embeddings to 1024 delivers better results on SCaR, we stick to selecting 256 as our configuration to balance computation and effectiveness.

### v2

@@ Abstract
Multimodal representation learning models have demonstrated successful operation across complex tasks, and the integration of vision-language models (VLMs) has further enabled embedding models with instruction-following capabilities. However, existing embedding models lack visual-interactive capabilities to specify regions of interest from users (e.g., point, bounding box, mask), which have been explored in generative models to broaden their human-interactive applicability. Equipping embedding models with visual interactions not only would unlock new applications with localized grounding of user intent, which remains unexplored, but also enable the models to learn entity-level information within images to complement their global representations for conventional embedding tasks. In this paper, we propose a novel Visual-InteRactive Text-Image Universal Embedder (VIRTUE) that extends the capabilities of the segmentation model and the vision-language model to the realm of representation learning. In VIRTUE, the segmentation model can process visual prompts that pinpoint specific regions within an image, thereby enabling the embedder to handle complex and ambiguous scenarios more precisely. To evaluate the visual-interaction ability of VIRTUE, we introduce a large-scale Segmentation-and-Scene Caption Retrieval (SCaR) benchmark comprising 1M samples that aims to retrieve the text caption by jointly considering the entity with a specific object and image scene. VIRTUE consistently achieves a state-of-the-art performance with significant improvements across 36 universal MMEB (3.1%–8.5%) and five visual-interactive SCaR (15.2%–20.3%) tasks. The code, models, and benchmarks are available at https://github.com/sony/virtue.
@@ 1 Introduction
Although VLM-based embedding models support interactive use, they rely on text as the primary human-machine interaction modality. In contrast, visual prompts, which have recently attracted attention in generative applications (You et al., 2024; Yuan et al., 2024; Lian et al., 2025), serve as an important but overlooked interaction channel. Visual prompting can not only enhance the downstream generation performance (Li et al., 2022b) but also provide precise spatial localization for fine-grained understanding (Liu et al., 2024c). This is particularly advantageous for embedding-based tasks, as it allows models to respond to visual inputs from the user beyond traditional global matching by capturing entity-level cues, thereby improving retrieval precision and alignment while complementing global representations.
@@
Method Novelty: We propose VIRTUE, a visual-interactive text-image universal embedder consisting of a VLM and a segmentation model to enable the visual interaction modality for human interactions. The segmentation model allows users to optionally provide different types of visual prompts via its prompt encoder and reinforces VIRTUE to capture entity-level representations in addition to global context. We also conduct a systematic analysis of visual prompt integration to provide guidelines for enabling visual-interactive capabilities in embedding models.
@@ 2 Related Works
Multimodal Representation Learning. Early progress in text-image representation learning was driven by two-tower contrastive models that learn a joint embedding space by aligning an image encoder with a text encoder (e.g., CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), SigLIP (Zhai et al., 2023), OpenCLIP (Cherti et al., 2023)). These models provide effective global image-text matching and have served as foundation models for building vision-language models (VLMs) that tackle zero-shot downstream tasks (Tong et al., 2024; Wang et al., 2025b; Deitke et al., 2025). Subsequent work has advanced embedding performance along various dimensions. For instance, UniIR (Wei et al., 2024) finetunes CLIP/BLIP with the late fusion of text and image embeddings, and UniME (Gu et al., 2025) distills text knowledge from an LLM followed by two-stage negatives contrastive learning. Magiclens (Zhang et al., 2024) incorporates open-ended instructions into dual-encoder architectures with training on large-scale instruction datasets. More recently, VLMs that accept arbitrary mixtures of visual and textual inputs and are trained with instruction-style objectives have emerged as flexible and unified embedding providers that better perform on multimodal reasoning and compositional queries, e.g., E5-V (Jiang et al., 2024), VLM2Vec (Jiang et al., 2025b), GME (Zhang et al., 2025), and LamRA (Liu et al., 2025b). Despite these advancements, existing embedding models only support textual instructions and lack native support for direct visual prompts; they are typically trained only on holistic image-text alignment (Oh et al., 2024). In contrast, VIRTUE integrates a segmentation model with a pretrained VLM to fuse segmentation-derived, prompt-conditioned entity representations with global representations, producing unified embeddings that are both entity-aware and context-preserving, as evaluated on both MMEB and our newly constructed SCaR.
@@ 3.1 SCaR Overview
Task Definition. The main difference between conventional image-to-text retrieval (e.g., MSCOCO_i2t in MMEB) and visual-interactive image-to-text retrieval is the additional region-of-interest input111Both retrieval scenarios include the text instruction (the MMEB instruction is shown in Tab. 10 of Jiang et al. (2025b)). SCaR uses the following instruction: “Find the caption that best describes the segmented object, considering both local details and global context in the given image.”. For simplicity, we omit these instructions here.. Formally, given an input image and a bounding box with ten candidates , the goal is to find the most relevant text caption , where is the embedding model and sim denotes cosine similarity. Since SCaR requires models to reason about both the specified object and the broader scene, candidate captions are intentionally challenging and demand joint reasoning over fine-grained bounding-box details and global image context.
@@ 3.2 Collection Pipeline
Fig. 2 illustrates the collection pipeline to construct SCaR, where all datasets are first converted to the COCO format to enable unified processing. To balance evaluation efficiency and coverage, we randomly sample up to five objects per image. For each sample, the image size, original caption, object category, and bounding box222While some datasets provide segmentation masks, we observed that GPT-4V struggles to interpret them reliably. In contrast, bounding boxes are easier for GPT-4V to parse and align with prior prompt-based collection strategies (Wang et al., 2025c; Jiang et al., 2025a); therefore, we opt for bounding boxes for SCaR. are provided in the prompt template, as shown in Fig. 5. The prompt instructs GPT-4V to return a JSON object containing the ground-truth caption and nine negatives. Since some datasets lack complete descriptions in terms of <object> <relation> <scene> (e.g., ADE20k only provides object names, and some RefCOCO+ samples do not contain global scene context), the prompt asks GPT-4V to verify whether the caption contains all three elements, and to supplement any missing elements with careful and image-grounded descriptions.
@@ 5.3 Comparison on SCaR for Visual-Interactive Tasks
5.4 Impacts of Various Visual Prompts on VIRTUE
To investigate the impacts of various visual prompts on VIRTUE, we analyze the robustness of VIRTUE-7B and VIRTUE-7B-SCaR-train under noisy and misaligned bounding boxes from the SCaR evaluation set. We further perform SCaR evaluations by replacing ground-truth bounding boxes with centroid points or random boxes, examining whether VIRTUE can effectively handle different types of interactive visual prompts for embedding generation. Finally, on MMEB, we study the effect of removing uniformly sampled points, either during inference only (w/o inference) or during both training and inference (w/o all), to understand how reliance on uniform point prompts influences performance.
| VIRTUE-7B | VIRTUE-7B-SCaR-train | |||||
| bbox | random | points | bbox | random | points | |
| SCaR | 27.8 | 9.2 | 26.6 | 56.9 | 13.0 | 45.0 |
| VIRTUE-7B | VIRTUE-7B-SCaR-train | |||||
| uniform | w/o inference | w/o all | uniform | w/o inference | w/o all | |
| MMEB | 68.6 | 24.1 | 65.5 | 66.8 | 41.9 | 43.4 |
Robustness to Noisy or Misaligned Visual Prompts. Fig. 5 presents results with different fractions of jittered bounding boxes, varying mask ratios for partial boxes, and off-by-k pixels. Both VIRTUE-7B and VIRTUE-7B-SCaR-train remain stable across different amounts of jitter and off-by-k offsets, demonstrating robustness of the segmentation streamline. On the other hand, using only 20% partial masks results for VIRTUE-7B-SCaR-train remains superior to baselines like VLM2Vec-7B-SCaR-train, suggesting that our model remains robust even when visual prompts provide limited coverage.
Effect of Using Points vs. (Random) Bounding Boxes. To understand if VIRTUE takes advantage of segmentation tokens, we hypothesize that if the model effectively leverages segmentation tokens, randomly sampling visual prompts should degrade performance; if it does not, performance should remain unchanged or improve (Makansi et al., 2022; Wang et al., 2025a). Tab. 5 shows the SCaR performance when bounding boxes (bbox), randomly sampled boxes (random), or centroid points (points) are adopted as visual prompts. Using points achieves performance comparable to bounding boxes for VIRTUE-7B, while VIRTUE-7B-SCaR-train exhibits a larger drop, indicating that finetuning on SCaR encourages a stronger reliance on bounding box prompts. The severe degradation with random boxes for both models further confirms that VIRTUE meaningfully uses the provided visual prompts when generating embeddings.
Impact of Removing Uniformly Sampled Points. Since VIRTUE incorporates uniformly sampled points as visual prompts for non-visual-interactive scenarios, Tab. 5 reports results obtained by removing these points either only during inference (w/o inference) or during both training and inference (w/o all). The performance declines in both settings indicate that uniform points help the segmentation streamline capture entity-level structure, which complements global context and leads to more fine-grained embeddings.
@@ 6 Conclusion
Acknowledgements
We gratefully acknowledge the support of all those who contributed to this research. We thank Naofumi Akimoto and Toshimitsu Uesaka for their valuable feedback and insightful suggestions that helped refine this work.
@@ Acknowledgements
The codebase, VIRTUE models, and SCaR benchmark samples are provided at https://github.com/sony/virtue to advance the community with the realm of interactive representation learning, and the experimental details as well as model configurations are summarized in Sec. 5.1 and Appendix C. Evaluations on the MMEB benchmark follows their official use, while SCaR is detailed in Sec. 3 and Appendix D.
@@ D.4 Additional SCaR Statistics
| Cohen’s Kappa | Drop samples (both annotators flagged) | Drop samples (either annotator flagged) |
| 0.89 | 4021 (7.7%) | 1638 (3.1%) |
We further analyze the composition of SCaR to better understand its benchmark characteristics. Tab. 7 presents the inter-annotator agreement and the number of dropped samples in the SCaR evaluation set. The high agreement score indicates that the two independent annotators were highly consistent in the dataset inspection process. After inspection, we removed approximately 11% of the samples: 7.7% were flagged by both annotators, and an additional 3.1% were flagged by either annotator.
Fig. 7 summarizes detailed statistics including word clouds, annotation counts, and image distributions for both the training and evaluation splits. Attributed to the creative and diverse instruction collection process described in Sec. 3.2, we observe that the word frequency distributions of ground-truth and negative captions are closely aligned across splits, suggesting that the benchmark is not dominated by a small set of frequent words. Moreover, both image- and annotation-level distributions reveal that COCO-Stuff and VisualGenome contribute the largest proportions of samples, reflecting their broad coverage of scenes and object relationships. Fig. 8 illustrates the sentence length distributions of ground-truth and negative captions across the five datasets in SCaR. Overall, the two distributions are closely aligned, indicating that the synthesized negative captions generated by GPT-4V preserve similar linguistic complexity to the ground-truth annotations rather than introducing trivial artifacts. We also observe slightly heavier tails in the negative caption distributions, suggesting greater variability in sentence length. This property enhances the linguistic diversity within SCaR while maintaining a comparable level of difficulty, thereby preventing models from exploiting superficial cues such as caption length.
@@ D.5 SCaR Examples
| Dataset | Query Image with bbox | Candidates |
@@
Motorcycle in forefront fully shown. Motorcycle in the garage fully shown. Motorcycle on a racetrack fully shown. Motorcycle in a field fully shown. Motorcycle with a helmet placed on the seat in forefront. Motorcycle being washed in forefront. Motorcycle loaded with packages in forefront. Bicycle in forefront fully shown. Scooter in forefront fully shown. Horse in forefront fully shown. |
@@
The bench closest to the palm tree and on a concrete pedestal at the beach. The bench closest to the palm tree and on a concrete pedestal in a city bus station. The bench closest to the palm tree and on a concrete pedestal in a shopping mall atrium. The bench closest to the palm tree and on a concrete pedestal in a hospital waiting area. The bench with a row of flower pots on its seat and on a concrete pedestal at the beach. The bench covered in colorful graffiti and on a concrete pedestal at the beach. The bench holding a stack of books and on a concrete pedestal at the beach. The playground slide closest to the palm tree and on a concrete pedestal at the beach. The trash can closest to the palm tree and on a concrete pedestal at the beach. The bicycle rack closest to the palm tree and on a concrete pedestal at the beach. |
@@
Mouth of sculpture by the waterfront. Mouth of sculpture in a museum gallery. Mouth of sculpture in a lush garden. Mouth of sculpture on a mountaintop. Mouth of sculpture blowing smoke by the waterfront. Mouth of sculpture illuminated by spotlights by the waterfront. Mouth of sculpture eating an apple by the waterfront. Fin of sculpture by the waterfront. Ear of sculpture by the waterfront. Tail of sculpture by the waterfront. |
@@
Cat sitting in front of a computer screen. Cat sitting on a kitchen countertop. Cat sitting in a garden. Cat sitting on a window sill. Cat pawing at a coffee cup in front of a computer screen. Cat curled up sleeping in front of a computer screen. Cat playing with headphones in front of a computer screen. Rabbit sitting in front of a computer screen. Dog sitting in front of a computer screen. Parrot sitting in front of a computer screen. |
@@
Fan standing near the chairs in a glass-roofed lounge. Fan standing near the chairs on a subway platform. Fan standing near the chairs in a hospital waiting area. Fan standing near the chairs in a gymnasium. Fan blowing onto a group of potted plants in a glass-roofed lounge. Fan hanging from the ceiling above the chairs in a glass-roofed lounge. Fan surrounded by scattered magazines on the floor in a glass-roofed lounge. Sculpture standing near the chairs in a glass-roofed lounge. Lamp standing near the chairs in a glass-roofed lounge. Plant standing near the chairs in a glass-roofed lounge. |
@@
Appendix E Additional Experiments
@@ E.1 Ablation Study
| 8) Alternative | MMEB | SCaR |
| Cross-attention | 63.1 | 30.0 |
| Attention-pool | 55.9 | 28.1 |
| Conv-MLP | 64.8 | 30.4 |
| 9) Finetune | MMEB | SCaR |
| ✗ | 64.8 | 30.4 |
| ✓ | 63.5 | 25.8 |
We conduct seven ablation variants to investigate the relative contributions of different components: 1) VLM backbones, 2) with and without an instruction in the query, 3) input image resolution, 4) segmentation streamline alternatives, 5) different sizes of SAM-2, 6) different numbers of MLPs in the segmentation-language connector (Eq. 1), where one MLP projects from to , and three MLPs project from to 768, then to 1024, and finally to ; 7) segmentation embedding lengths ; 8) cross-modal modeling alternatives; and 9) finetuning and freezing SAM-2. For 4) segmentation streamline alternatives, we replace our segmentation model with Prompter from CLOC555CLOC cannot be directly evaluated or reproduced, as its model, code, and parts of training data are not publicly available. We attempted to contact the authors but did not receive a response. (Chen et al., 2025), which encodes bounding boxes via a randomly initialized Transformer, or directly crop images for the vision encoder (denoted as Cropped). Both variants concatenate embeddings from nine random crops, similar to the setting of nine sampled points in SAM2.1 for the segmentation-language connector for MMEB. For 8) cross-modal modeling alternatives, we adopt two variants for the segmentation-language connect: Cross-attention, which uses the same architecture as the vision-language connector in Qwen2-VL. Attention-pooling: A natural approach using learnable queries to pool information from the region features (we used ).
As shown in Tab. 9, although Phi-3.5-V has more parameters than Qwen2-VL-2B, the latter yields better results on MMEB and SCaR. Removing text instructions and using lower resolutions lead to a significantly inferior performance. The degraded results from Prompter and cropping suggest that pre-trained SAM2.1 provides not only more precise segmentations than simple crops but also a stronger understanding from its pre-trained knowledge. Although SAM2.1-L performs on par with SAM2.1-B+ in terms of MMEB, its performance degrades on SCaR. In contrast, using SAM2.1-S deleteriously impacts both benchmarks. Additionally, increasing MLP layers and segmentation embedding length does not consistently yield gains across MMEB and SCaR. While increasing segmentation embeddings to 1024 delivers better results on SCaR, we stick to selecting 256 as our configuration to balance computation and effectiveness. Using attention-based cross-modal methods does not yield better results, although cross-attention achieves performance comparable to our convolutional MLPs. Similarly, finetuning SAM-2 performs worse than keeping it frozen–especially on SCaR–since MMEB-train does not contain user-specific visual prompts (i.e., using uniform points). This mismatch biases the segmentation pipeline toward non–visual-interactive scenarios.